OpenStudy (anonymous):

Find the limit of the function by using direct substitution. limit as x approaches three of quantity x squared plus eight x minus two.

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3}(x^{2} + 8x - 2)\]

Nnesha (nnesha):

\[\large\rm \lim_{x \rightarrow 3} ~~~x^2+8x-2\] like this ??

Nnesha (nnesha):

ohh okay cool. *direct substitution* just substitute x for 3 and then solve. that's it.

OpenStudy (anonymous):

\[\lim_{x \rightarrow \frac{ \pi }{ 2 }}(2e ^{x}\sin x)\]

OpenStudy (anonymous):

i have to use direct substitution for this one as well

Nnesha (nnesha):

alright same steps when the direction is *direct substitution * then just substitute x for the number its approaches by

OpenStudy (anonymous):

\[(2e ^{\left(\begin{matrix}\pi \\ 2\end{matrix}\right)} \sin \frac{ \pi }{ 2 })\]

OpenStudy (anonymous):

how do i solve from there?

Nnesha (nnesha):

sin (pi/2) = ??? you can use unit circle.

OpenStudy (anonymous):

equals 0